# Existence of non-Abelian vortices in a coupled 4D–2D quantum field theory

### v1

@@ 2 Gauged non–Abelian vortex model
For a minimal vortex with a fixed orientation in color–flavor, for example $(1,\mathbf{1}_{N-1})$, we can choose the spherically–symmetric ansatz in the scalar field
@@
For the nonlinear ordinary differential equations (2.18)–(2.21) subject to the boundary conditions (2.22)–(2.23), we know that the existence of solutions can be obtained by numerical methods in literature [7].
@@ 3 Governing system of nonlinear elliptic equations
$rQ^{\prime}_{2}-Q_{2}\left(\frac{-f_{NA}+f}{N}\right)$ $=$ 00 0 (3.4)
and the boundary conditions
$f(r)$ $=$ $1,~{}~{}~{}f_{NA}(r)=1,~{}~{}~{}\text{as}~{}~{}r\to 0,$ (3.5)
$f(r)$ $=$ $0,~{}~{}~{}f_{NA}(r)=0,~{}~{}~{}Q_{1}(r)=1,~{}~{}~{}Q_{2}(r)=1,~{}~{}~{}\text{as}~{}~{}r\to\infty.$ (3.6)
It is worth noting that $Q_{1}(r)\neq 0$ for all $r\in(0,+\infty)$. If otherwise, $Q_{1}(r)=0$ at some $r\in(0,+\infty)$, then $Q_{1}(r)\equiv 0$ for all $r\in(0,+\infty)$ by the continuous dependence theorem for solutions of the initial value problems of ordinary differential equations. Similarly, $Q_{2}(r)\neq 0$ for all $r\in(0,+\infty)$. As a consequence, the equations (3.3)–(3.4) are recast into
@@
where $N>1$ is an integer and $u_{1}(+\infty)=u_{2}(+\infty)=0$. By a direct computation, we obtain the following system of nonlinear elliptic equations
$\triangle u_{1}(x)=\left(\frac{3}{2}-\frac{1}{2N}\right)\left(\mathrm{e}^{2u_{1}(x)}-1\right)+\left(N-\frac{3}{2}+\frac{1}{2N}\right)\left(\mathrm{e}^{2u_{2}(x)}-1\right),$ (3.11)
$\triangle u_{2}(x)=\left(1-\frac{1}{2N}\right)\left(\mathrm{e}^{2u_{1}(x)}-1\right)+\left(N-1+\frac{1}{2N}\right)\left(\mathrm{e}^{2u_{2}(x)}-1\right),$ (3.12)
where $N>1$ and $x\in\mathbb{R}^{2}$. We are interested in the existence of solutions of (3.11)–(3.12) for the case of the full plane. Therefore, we consider the system (3.11)–(3.12) over the plane with the topological boundary condition
@@
the equations (3.11)–(3.12) can be rewritten in a compact form
$\triangle u_{i}=\sum_{j=1}^{2}a_{ij}\left(\mathrm{e}^{2u_{j}}-1\right),~{}~{}~{}i=1,2.$ (3.15)
@@
$\triangle u_{1}=\frac{5}{4}\left(\mathrm{e}^{2u_{1}}-1\right)+\frac{3}{4}\left(\mathrm{e}^{2u_{2}}-1\right),$ (3.16)
$\triangle u_{2}=\frac{3}{4}\left(\mathrm{e}^{2u_{1}}-1\right)+\frac{5}{4}\left(\mathrm{e}^{2u_{2}}-1\right).$ (3.17)
@@
$\left|pu_{1}(x)\right|^{2}+\left|nu_{2}(x)\right|^{2}\leq C(\varepsilon)\mathrm{e}^{-(1-\varepsilon)\sqrt{\lambda_{0}}|x|},$ (3.18)
$\left|\nabla\left(mu_{1}(x)+nu_{2}(x)\right)\right|^{2}+\left|\nabla\left(pu_{1}(x)+qu_{2}(x)\right)\right|^{2}\leq C(\varepsilon)\mathrm{e}^{-(1-\varepsilon)\sqrt{\lambda}|x|},$ (3.19)
@@
$m=\frac{\left(2\alpha-1\right)^{2}}{2\beta\left(\lambda_{3}-\alpha\right)},~{}~{}n=2,~{}~{}p=\frac{2\alpha-1}{\beta},~{}~{}q=\frac{4\left(\lambda_{4}-\alpha\right)}{2\alpha-1},$ (3.20)
@@
$\int_{\mathbb{R}^{2}}\left\{\left[\left(m+n\right)\alpha-\frac{n}{2}\right]\left(\mathrm{e}^{2u_{1}(x)}-1\right)+\left[\left(m+n\right)\beta+\frac{n}{2}\right]\left(\mathrm{e}^{2u_{2}(x)}-1\right)\right\}\mathrm{d}x=0,$
$\int_{\mathbb{R}^{2}}\left\{\left[\left(p+q\right)\alpha-\frac{q}{2}\right]\left(\mathrm{e}^{2u_{1}(x)}-1\right)+\left[\left(p+q\right)\beta+\frac{q}{2}\right]\left(\mathrm{e}^{2u_{2}(x)}-1\right)\right\}\mathrm{d}x=0.$ (3.21)
@@ 4.1 Existence and uniqueness of the critical point
$\mathbf{u}=\left(u_{1},u_{2}\right)^{\tau},~{}~{}~{}\mathbf{E}=\left(\mathrm{e}^{2u_{1}}-1,\mathrm{e}^{2u_{2}}-1\right)^{\tau}$ (4.1)
@@
$\triangle\mathbf{u}=A\mathbf{E}.$ (4.2)
@@
$\mathbf{w}=L^{-1}\mathbf{u}~{}~{}~{}\text{or}~{}~{}~{}\mathbf{u}=L\mathbf{w},$ (4.6)
@@
$\triangle\mathbf{w}=R\mathbf{E}.$ (4.7)
@@
$\triangle w_{1}=\left(\frac{3}{2}-\frac{1}{2N}\right)\left(\mathrm{e}^{2w_{1}}-1\right)+\left(N-\frac{3}{2}+\frac{1}{2N}\right)\left(\mathrm{e}^{2\left(\frac{2N-1}{3N-1}w_{1}+w_{2}\right)}-1\right),$ (4.8)
$\triangle w_{2}=\frac{N^{2}}{3N-1}\left(\mathrm{e}^{2\left(\frac{2N-1}{3N-1}w_{1}+w_{2}\right)}-1\right).$ (4.9)
In order to see the variational structure of (4.8)–(4.9) clearly, it is beneficial to reformulate them equivalently as
$\frac{4N}{(3N-1)(N-1)}\triangle w_{1}=\frac{2}{N-1}\left(\mathrm{e}^{2w_{1}}-1\right)+\frac{2(2N-1)}{3N-1}\left(\mathrm{e}^{2\left(\frac{2N-1}{3N-1}w_{1}+w_{2}\right)}-1\right),$ (4.10)
$\frac{2(3N-1)}{N^{2}}\triangle w_{2}=2\left(\mathrm{e}^{2\left(\frac{2N-1}{3N-1}w_{1}+w_{2}\right)}-1\right).$ (4.11)
@@
$I(w_{1},w_{2})=\int_{\mathbb{R}^{2}}\Bigg{\{}\frac{2N}{(3N-1)(N-1)}\left|\nabla w_{1}\right|^{2}+\frac{3N-1}{N^{2}}\left|\nabla w_{2}\right|^{2}+\left(\mathrm{e}^{2\left(\frac{2N-1}{3N-1}w_{1}+w_{2}\right)}-1\right)$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\frac{1}{N-1}\left(\mathrm{e}^{2w_{1}}-1\right)-\frac{4N^{2}}{(3N-1)(N-1)}w_{1}-2w_{2}\Bigg{\}}\mathrm{d}x.$ (4.12)
@@
$DI(w_{1},w_{2})(w_{1},w_{2})=\int_{\mathbb{R}^{2}}\Bigg{\{}\frac{4N}{(3N-1)(N-1)}\left|\nabla w_{1}\right|^{2}+\frac{6N-2}{N^{2}}\left|\nabla w_{2}\right|^{2}-\frac{4N^{2}}{(3N-1)(N-1)}w_{1}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-2w_{2}+\frac{2w_{1}}{N-1}\mathrm{e}^{2w_{1}}+2\left(\frac{2N-1}{3N-1}w_{1}+w_{2}\right)\mathrm{e}^{2\left(\frac{2N-1}{3N-1}w_{1}+w_{2}\right)}\Bigg{\}}\mathrm{d}x$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}=\int_{\mathbb{R}^{2}}\Bigg{\{}\frac{4N}{(3N-1)(N-1)}\left|\nabla w_{1}\right|^{2}+\frac{6N-2}{N^{2}}\left|\nabla w_{2}\right|^{2}+\frac{2w_{1}}{N-1}\left(\mathrm{e}^{2w_{1}}-1\right)$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+2\left(\frac{2N-1}{3N-1}w_{1}+w_{2}\right)\left(\mathrm{e}^{2\left(\frac{2N-1}{3N-1}w_{1}+w_{2}\right)}-1\right)\Bigg{\}}\mathrm{d}x.$ (4.13)
@@
$\left|\nabla w_{1}\right|^{2}+\left|\nabla w_{2}\right|^{2}=\left|\nabla u_{1}\right|^{2}+\left|\nabla\left(u_{2}-\gamma u_{1}\right)\right|^{2}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\,\leq\left|\nabla u_{1}\right|^{2}+\left|\nabla u_{2}\right|^{2}+\gamma^{2}\left|\nabla u_{1}\right|^{2}+2\gamma\left|\left(\nabla u_{1},\nabla u_{2}\right)\right|$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\,\leq\left(1+\gamma+\gamma^{2}\right)\left|\nabla u_{1}\right|^{2}+\left(1+\gamma\right)\left|\nabla u_{2}\right|^{2}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\,<3\left(\left|\nabla u_{1}\right|^{2}+\left|\nabla u_{2}\right|^{2}\right),$ (4.14)
@@
$\left|\nabla w_{1}\right|^{2}+\left|\nabla w_{2}\right|^{2}\geq\left(1+\gamma^{2}\right)\left|\nabla u_{1}\right|^{2}+\left|\nabla u_{2}\right|^{2}-2\gamma\left|\left(\nabla u_{1},\nabla u_{2}\right)\right|$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\,\geq\left(1+\gamma^{2}-\frac{\gamma}{2\varepsilon}\right)\left|\nabla u_{1}\right|^{2}+\left(1-2\gamma\varepsilon\right)\left|\nabla u_{2}\right|^{2}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\,\geq\left(\frac{5}{4}-\frac{1}{3\varepsilon}\right)\left|\nabla u_{1}\right|^{2}+\left(1-\frac{4\varepsilon}{3}\right)\left|\nabla u_{2}\right|^{2},$ (4.15)
@@
$\left|\nabla w_{1}\right|^{2}+\left|\nabla w_{2}\right|^{2}\geq\frac{1}{4}\left|\nabla u_{1}\right|^{2}+\frac{5}{9}\left|\nabla u_{2}\right|^{2}>\frac{5}{9}\left(\left|\nabla u_{1}\right|^{2}+\left|\nabla u_{2}\right|^{2}\right).$ (4.16)
@@
$\frac{5}{9}\left(\left|\nabla u_{1}\right|^{2}+\left|\nabla u_{2}\right|^{2}\right)<\left|\nabla w_{1}\right|^{2}+\left|\nabla w_{2}\right|^{2}<3\left(\left|\nabla u_{1}\right|^{2}+\left|\nabla u_{2}\right|^{2}\right).$ (4.17)
@@
$\frac{5}{9}\left(u_{1}^{2}+u_{2}^{2}\right)<w_{1}^{2}+w_{2}^{2}<3\left(u_{1}^{2}+u_{2}^{2}\right).$ (4.18)
@@
$DI(w_{1},w_{2})(w_{1},w_{2})-\frac{20N}{9(3N-1)(N-1)}\int_{\mathbb{R}^{2}}\left(\left|\nabla u_{1}\right|^{2}+\left|\nabla u_{2}\right|^{2}\right)\mathrm{d}x$
@@
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\geq\int_{\mathbb{R}^{2}}\Bigg{\{}\frac{2u_{1}}{N-1}\left(\mathrm{e}^{2u_{1}}-1\right)+2u_{2}\left(\mathrm{e}^{2u_{2}}-1\right)\Bigg{\}}\mathrm{d}x$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\equiv M_{1}(u_{1})+M_{2}(u_{2}).$ (4.19)

### v2

@@ 2 Gauged non–Abelian vortex model
For a minimal vortex with a fixed orientation in color–flavor, for example $(1,\mathbf{1}_{N-1})$, we can choose the radially symmetric ansatz in the scalar field
@@
It should be noted that the finite vortex points of both $Q_{1}$ and $Q_{2}$ are concentrated at the origin in the radial symmetry situation, that is to say, $Q_{1}(r)=Q_{2}(r)=0$ as $r\to 0$. Moreover, in view of the equations (2.18)–(2.21), we see that $f$ is not equal to $f_{NA}$ and $Q_{1}$ is not equal to $Q_{2}$. For the nonlinear ordinary differential equations (2.18)–(2.21) subject to the boundary conditions (2.22)–(2.23), we know that the existence of solutions can be obtained by numerical methods in literature [7].
@@ 3 Governing system of nonlinear elliptic equations
$rQ^{\prime}_{2}-Q_{2}\left(\frac{-f_{NA}+f}{N}\right)$ $=$ $0$ (3.4)
It is worth noting that $Q_{1}(r)\neq 0$ for all $r\in(0,+\infty)$. If otherwise, $Q_{1}(r)=0$ at some $r\in(0,+\infty)$, then $Q_{1}(r)\equiv 0$ for all $r\in(0,+\infty)$ by the continuous dependence theorem for solutions of the initial value problems of ordinary differential equations. Similarly, $Q_{2}(r)\neq 0$ for all $r\in(0,+\infty)$. As a consequence, when $r>0$ the equations (3.3)–(3.4) are recast into
@@
where $N>1$ is an integer. By a direct computation, we obtain the following system of nonlinear elliptic equations over $\mathbb{R}^{2}$
$\triangle u_{1}(x)$ $=$ $\left(\frac{3}{2}-\frac{1}{2N}\right)\left(\mathrm{e}^{2u_{1}(x)}-1\right)+\left(N-\frac{3}{2}+\frac{1}{2N}\right)\left(\mathrm{e}^{2u_{2}(x)}-1\right)+4\pi n_{1}\delta(x),$ (3.9)
$\triangle u_{2}(x)$ $=$ $\left(1-\frac{1}{2N}\right)\left(\mathrm{e}^{2u_{1}(x)}-1\right)+\left(N-1+\frac{1}{2N}\right)\left(\mathrm{e}^{2u_{2}(x)}-1\right)+4\pi n_{2}\delta(x),$ (3.10)
where $\delta(x)$ is the Dirac distribution on $\mathbb{R}^{2}$ concentrated at the origin and the positive integers $n_{1}$, $n_{2}$ are the multiplicities of the vortices corresponding to $u_{1}$ and $u_{2}$. We are interested in the existence of solutions of (3.9)–(3.10) for the case of the full plane. Therefore, we consider the system (3.9)–(3.10) over the plane with the topological boundary condition
@@
then the equations (3.9)–(3.10) can be rewritten in a compact form
$\triangle u_{i}=\sum_{j=1}^{2}a_{ij}\left(\mathrm{e}^{2u_{j}}-1\right)+4\pi n_{i}\delta(x),~{}~{}~{}i=1,2.$ (3.13)
@@
$\triangle u_{1}$ $=$ $\frac{5}{4}\left(\mathrm{e}^{2u_{1}}-1\right)+\frac{3}{4}\left(\mathrm{e}^{2u_{2}}-1\right)+4\pi n_{1}\delta(x),$ (3.14)
$\triangle u_{2}$ $=$ $\frac{3}{4}\left(\mathrm{e}^{2u_{1}}-1\right)+\frac{5}{4}\left(\mathrm{e}^{2u_{2}}-1\right)+4\pi n_{2}\delta(x).$ (3.15)
@@
$\left|pu_{1}(x)\right|^{2}+\left|2u_{2}(x)\right|^{2}\leq C(\varepsilon)\mathrm{e}^{-(1-\varepsilon)\sqrt{\lambda_{0}}|x|},$ (3.16)
$\left|\nabla\left(mu_{1}(x)+2u_{2}(x)\right)\right|^{2}+\left|\nabla\left(pu_{1}(x)+qu_{2}(x)\right)\right|^{2}\leq C(\varepsilon)\mathrm{e}^{-(1-\varepsilon)\sqrt{\lambda}|x|},$ (3.17)
@@
$m=\frac{\left(2\alpha-1\right)^{2}}{2\beta\left(\lambda_{3}-\alpha\right)},~{}~{}p=\frac{2\alpha-1}{\beta},~{}~{}q=\frac{4\left(\lambda_{4}-\alpha\right)}{2\alpha-1},$ (3.18)
@@
$\int_{\mathbb{R}^{2}}\left\{\left[\left(m+2\right)\alpha-1\right]\left(\mathrm{e}^{2u_{1}(x)}-1\right)+\left[\left(m+2\right)\beta+1\right]\left(\mathrm{e}^{2u_{2}(x)}-1\right)\right\}\mathrm{d}x=-4\pi\left(mn_{1}+2n_{2}\right),$
$\int_{\mathbb{R}^{2}}\left\{\left[\left(p+q\right)\alpha-\frac{q}{2}\right]\left(\mathrm{e}^{2u_{1}(x)}-1\right)+\left[\left(p+q\right)\beta+\frac{q}{2}\right]\left(\mathrm{e}^{2u_{2}(x)}-1\right)\right\}\mathrm{d}x=-4\pi\left(pn_{1}+qn_{2}\right),$
where $\alpha$, $\beta$ are shown in (3.12) and $n_{1}$, $n_{2}$ are the multiplicities of the vortices corresponding to $u_{1}$ and $u_{2}$.
@@ 4.1 Existence and uniqueness of the critical point
As in [16], we need to take the background function
$u_{i}^{0}(x)=-n_{i}\mathrm{ln}\left(1+\tau|x|^{-2}\right),~{}~{}~{}i=1,2,~{}~{}~{}\tau>0,$ (4.1)
then we can see that
$\triangle u_{i}^{0}(x)=-\varphi_{i}(x)+4\pi n_{i}\delta(x),~{}~{}~{}\varphi_{i}(x)=\frac{4n_{i}\tau}{\left(\tau+|x|^{2}\right)^{2}},~{}~{}~{}i=1,2.$ (4.2)
It is significant noting that $\int_{\mathbb{R}^{2}}\varphi_{i}(x)\mathrm{d}x=4\pi n_{i}(i=1,2)$. Setting $u_{i}(x)=u_{i}^{0}(x)+P_{i}(x)(i=1,2)$, the system (3.13) becomes
[…]
@@
$\mathbf{P}=\left(P_{1},P_{2}\right)^{\tau},~{}~{}~{}\mathbf{E}=\left(\mathrm{e}^{2(u_{1}^{0}+P_{1})}-1,\mathrm{e}^{2(u_{2}^{0}+P_{2})}-1\right)^{\tau},~{}~{}~{}\mathbf{\Phi}=\left(\varphi_{1},\varphi_{2}\right)^{\tau}$ (4.5)
@@
$\triangle\mathbf{P}=A\mathbf{E}+\mathbf{\Phi}.$ (4.6)
@@
$\mathbf{w}=L^{-1}\mathbf{P}~{}~{}~{}\text{or}~{}~{}~{}\mathbf{P}=L\mathbf{w},$ (4.10)
@@
$\triangle\mathbf{w}=R\mathbf{E}+L^{-1}\mathbf{\Phi}.$ (4.11)
@@
$\triangle w_{1}$ $=$ $\alpha\left(\mathrm{e}^{2\left(u_{1}^{0}+w_{1}\right)}-1\right)+\beta\left(\mathrm{e}^{2\left[u_{2}^{0}+\left(1-\frac{1}{2\alpha}\right)w_{1}+w_{2}\right]}-1\right)+\psi_{1},$ (4.12)
$\triangle w_{2}$ $=$ $\frac{\alpha+\beta}{2\alpha}\left(\mathrm{e}^{2\left[u_{2}^{0}+\left(1-\frac{1}{2\alpha}\right)w_{1}+w_{2}\right]}-1\right)+\psi_{2},$ (4.13)
where $\psi_{1}=\varphi_{1}$, $\psi_{2}=(\frac{1}{2\alpha}-1)\varphi_{1}+\varphi_{2}$. In order to see the variational structure of (4.12)–(4.13) clearly, it is beneficial to reformulate them equivalently as
$\frac{2\alpha-1}{\alpha\beta}\triangle w_{1}$ $=$ $\frac{2\alpha-1}{\beta}\left(\mathrm{e}^{2\left(u_{1}^{0}+w_{1}\right)}-1\right)+\left(2-\frac{1}{\alpha}\right)\left(\mathrm{e}^{2\left[u_{2}^{0}+\left(1-\frac{1}{2\alpha}\right)w_{1}+w_{2}\right]}-1\right)$ (4.14)
$+$ $\frac{2\alpha-1}{\alpha\beta}\psi_{1},$
$\frac{4\alpha}{\alpha+\beta}\triangle w_{2}$ $=$ $2\left(\mathrm{e}^{2\left[u_{2}^{0}+\left(1-\frac{1}{2\alpha}\right)w_{1}+w_{2}\right]}-1\right)+\frac{4\alpha}{\alpha+\beta}\psi_{2}.$ (4.15)
@@
$I(w_{1},w_{2})=\int_{\mathbb{R}^{2}}\Bigg{\{}\frac{2\alpha-1}{2\alpha\beta}\left|\nabla w_{1}\right|^{2}+\frac{2\alpha}{\alpha+\beta}\left|\nabla w_{2}\right|^{2}+\mathrm{e}^{2u_{2}^{0}}\left(\mathrm{e}^{2\left[\left(1-\frac{1}{2\alpha}\right)w_{1}+w_{2}\right]}-1\right)$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\frac{2\alpha-1}{2\beta}\mathrm{e}^{2u_{1}^{0}}\left(\mathrm{e}^{2w_{1}}-1\right)+\frac{2\alpha-1}{\alpha\beta}\psi_{1}w_{1}-\frac{\left(2\alpha-1\right)\left(\alpha+\beta\right)}{\alpha\beta}w_{1}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\frac{4\alpha}{\alpha+\beta}\psi_{2}w_{2}-2w_{2}\Bigg{\}}\mathrm{d}x.$ (4.16)
@@
$DI(w_{1},w_{2})(w_{1},w_{2})=\int_{\mathbb{R}^{2}}\Bigg{\{}\frac{2\alpha-1}{\alpha\beta}\left|\nabla w_{1}\right|^{2}+\frac{4\alpha}{\alpha+\beta}\left|\nabla w_{2}\right|^{2}+\left(\mathrm{e}^{2u_{2}^{0}}-1\right)\left[\left(2-\frac{1}{\alpha}\right)w_{1}+2w_{2}\right]$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\frac{2\alpha-1}{\beta}\left(\mathrm{e}^{2u_{1}^{0}}-1\right)w_{1}+\frac{2\alpha-1}{\beta}\mathrm{e}^{2u_{1}^{0}}\left(\mathrm{e}^{2w_{1}}-1\right)w_{1}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\mathrm{e}^{2u_{2}^{0}}\left(\mathrm{e}^{\left(2-\frac{1}{\alpha}\right)w_{1}+2w_{2}}-1\right)\left[\left(2-\frac{1}{\alpha}\right)w_{1}+2w_{2}\right]$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\frac{2\alpha-1}{\alpha\beta}\psi_{1}w_{1}+\frac{4\alpha}{\alpha+\beta}\psi_{2}w_{2}\Bigg{\}}\mathrm{d}x.$ (4.17)
@@
$\left|\nabla w_{1}\right|^{2}+\left|\nabla w_{2}\right|^{2}=\left|\nabla P_{1}\right|^{2}+\left|\nabla\left(P_{2}-\gamma P_{1}\right)\right|^{2}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\,\leq\left|\nabla P_{1}\right|^{2}+\left|\nabla P_{2}\right|^{2}+\gamma^{2}\left|\nabla P_{1}\right|^{2}+2\gamma\left|\left(\nabla P_{1},\nabla P_{2}\right)\right|$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\,\leq\left(1+\gamma+\gamma^{2}\right)\left|\nabla P_{1}\right|^{2}+\left(1+\gamma\right)\left|\nabla P_{2}\right|^{2}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\,<3\left(\left|\nabla P_{1}\right|^{2}+\left|\nabla P_{2}\right|^{2}\right),$ (4.18)
@@
$\left|\nabla w_{1}\right|^{2}+\left|\nabla w_{2}\right|^{2}\geq\left(1+\gamma^{2}\right)\left|\nabla P_{1}\right|^{2}+\left|\nabla P_{2}\right|^{2}-2\gamma\left|\left(\nabla P_{1},\nabla P_{2}\right)\right|$ (4.19)
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\,\geq\left(1+\gamma^{2}-\frac{\gamma}{2\varepsilon}\right)\left|\nabla P_{1}\right|^{2}+\left(1-2\gamma\varepsilon\right)\left|\nabla P_{2}\right|^{2}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\,\geq\left(\frac{5}{4}-\frac{1}{3\varepsilon}\right)\left|\nabla P_{1}\right|^{2}+\left(1-\frac{4\varepsilon}{3}\right)\left|\nabla P_{2}\right|^{2},$
@@
$\left|\nabla w_{1}\right|^{2}+\left|\nabla w_{2}\right|^{2}\geq\frac{1}{4}\left|\nabla P_{1}\right|^{2}+\frac{5}{9}\left|\nabla P_{2}\right|^{2}>\frac{5}{9}\left(\left|\nabla P_{1}\right|^{2}+\left|\nabla P_{2}\right|^{2}\right).$ (4.20)
@@
$\frac{5}{9}\left(\left|\nabla P_{1}\right|^{2}+\left|\nabla P_{2}\right|^{2}\right)<\left|\nabla w_{1}\right|^{2}+\left|\nabla w_{2}\right|^{2}<3\left(\left|\nabla P_{1}\right|^{2}+\left|\nabla P_{2}\right|^{2}\right).$ (4.21)
@@
$\frac{5}{9}\left(P_{1}^{2}+P_{2}^{2}\right)<w_{1}^{2}+w_{2}^{2}<3\left(P_{1}^{2}+P_{2}^{2}\right).$ (4.22)
@@
$DI(w_{1},w_{2})(w_{1},w_{2})-\frac{5}{9}\min\left\{\frac{2\alpha-1}{2\beta},\frac{4\alpha}{\alpha+\beta}\right\}\int_{\mathbb{R}^{2}}\left(\left|\nabla P_{1}\right|^{2}+\left|\nabla P_{2}\right|^{2}\right)\mathrm{d}x$
@@
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\geq\int_{\mathbb{R}^{2}}\Bigg{\{}\frac{2\alpha-1}{\beta}\left(\mathrm{e}^{2(u_{1}^{0}+P_{1})}-1\right)P_{1}+2\left[\frac{2\alpha-1}{2\alpha\beta}\psi_{1}+\frac{1-2\alpha}{\alpha+\beta}\psi_{2}\right]P_{1}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+2\left(\mathrm{e}^{2(u_{2}^{0}+P_{2})}-1\right)P_{2}+\frac{4\alpha}{\alpha+\beta}\psi_{2}P_{2}\Bigg{\}}\mathrm{d}x$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}=\frac{2\alpha-1}{2\beta}\int_{\mathbb{R}^{2}}2P_{1}\left(\mathrm{e}^{2(u_{1}^{0}+P_{1})}-1+X_{1}\right)\mathrm{d}x+\int_{\mathbb{R}^{2}}2P_{2}\left(\mathrm{e}^{2(u_{2}^{0}+P_{2})}-1+X_{2}\right)\mathrm{d}x$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\equiv\frac{2\alpha-1}{2\beta}M_{1}(P_{1})+M_{2}(P_{2}),$ (4.23)
where
$X_{1}=\frac{1}{\alpha}\psi_{1}-\frac{2\beta}{\alpha+\beta}\psi_{2},~{}~{}~{}X_{2}=\frac{2\alpha}{\alpha+\beta}\psi_{2}.\$